# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### v2

@@ Introduction
The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II. Since the only theory paper (Bian et al. 2023) on the SMS-EMOA merely considers its performance on the bi-objective OJZJ problem, to broaden our understanding of this algorithm we now analyze its runtime for the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
OneMinMax and LOTZ
@@
The OneMinMax benchmark introduced by Giel and Lehre (2010) (and also the CountingOnesCountingZeroes benchmark proposed by Laumanns et al. (2002)) are bi-objective counterparts of the famous single-objective OneMax benchmark. Likewise, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler (2004) (also the weighted version WLPTNO of Qian, Yu, and Zhou (2013)) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
@@
For OneMinMax, the two objectives count the number of ones and zeros in a given bit-string, respectively. For LOTZ, the first objective is the number of contiguous ones starting from the first bit position, and the second objective is the number of contiguous zeros starting from the last bit position. Both benchmarks with their simple and clear structure facilitate the fundamental theoretical understanding of MOEAs. We note that there are other benchmarks used in the mathematical analysis of MOEAs (see, e.g. Horoba and Neumann 2008; Brockhoff et al. 2009; Qian, Tang, and Zhou 2016; Li et al. 2016; Dang et al. 2023b), but clearly OneMinMax and LOTZ are the most prominent ones.
Here are the formal definitions of OneMinMax and LOTZ.
Definition 18 (Giel and Lehre (2010)).
@@
Let $n\in\mathbb{N}$ be the problem size. The OneMinMax function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{OneMinMax}(x)=\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
Definition 19 (Laumanns, Thiele, and Zitzler (2004)).
Let $n\in\mathbb{N}$ be the problem size. The LOTZ function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{LOTZ}(x)=\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
Runtime
@@
We now analyze the runtime of the SMS-EMOA on the OneMinMax and LOTZ benchmarks. Note that the Pareto front, and more generally the size of any set of pairwise non-dominated solutions, for OneMinMax and LOTZ have sizes at most $n+1$. Hence we can use Lemma 4 with $M=n+1$.
@@
Theorem 20 below gives an upper bound for the runtime of SMS-EMOA on OneMinMax. This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022). Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our results consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
@@
Our runtime guarantee for LOTZ, see Theorem 21 below, again coincides with the guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022). In the proof, the process is divided into two stages. The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point. For the first stage, we pessimistically consider the time to reach $1^{n}$. For the second stage, we consider the time to generate $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other. In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost, which allows to add waiting times of successive progress events.
@@
Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations (or fitness evaluations) in expectation, the full Pareto front is covered.
Note that the runtimes of the GSEMO on OneMinMax and LOTZ are $O(n^{2}\ln n)$ and $O(n^{3})$, respectively (Giel and Lehre 2010; Laumanns, Thiele, and Zitzler 2004). Hence, they agree with our bounds for the SMS-EMOA when using the most interesting population size $\mu=\Theta(n)$.
Conclusion
@@
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr (2023) supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
@@
We also showed that the stochastic population update proposed in (Bian et al. 2023) for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.
@@
On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain when increasing the number of objectives.
@@
Given that our first result shows advantages of the SMS-EMOA and there is only one previous work performing rigorous runtime analyses for this algorithm, we extended our knowledge in this direction by proving competitive runtime guarantees for this algorithm on the two most prominent bi-objective benchmarks.

### v3

@@ 7 Runtime Analysis for Two Classic Bi-objective Benchmarks
In the sections above, we discussed the runtime of the SMS-EMOA for the many-objective OJZJ benchmark and showed a good performance, different from the original NSGA-II. Since the only two previous theory papers [BZLQ23, ZLDD24] on the SMS-EMOA merely considers its performance on the bi-objective OJZJ and DLTB problems, to broaden our understanding of this algorithm, we now analyze its runtime for the two most prominent bi-objective benchmarks, OneMinMax and LOTZ.
7.1 OneMinMax and LOTZ
@@
The OneMinMax benchmark introduced by Giel and Lehre [GL10] (and likewise the similar CountingOnesCountingZeroes benchmark previously proposed by Laumanns et al. [LTZ+02]) are bi-objective counterparts of the famous single-objective OneMax benchmark. Analogously, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler [LTZ04] (also the weighted version WLPTNO of Qian, Yu, and Zhou [QYZ13]) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
@@
For OneMinMax, the two objectives count the number of ones and zeros in a given bit-string, respectively. For LOTZ, the first objective is the number of contiguous ones starting from the first bit position, and the second objective is the number of contiguous zeros starting from the last bit position. Both benchmarks with their simple and clear structure have greatly facilitated the understanding of how MOEAs work. We note that there are other benchmarks used in the mathematical analysis of MOEAs, see, e.g., [HN08, BFH+09, QTZ16, LZZZ16, DOSS23b], but clearly OneMinMax and LOTZ are the most prominent ones.
Definition 25 ([GL10, LTZ04]).
@@
Let $n\in\mathbb{N}$ be the problem size. The OneMinMax and LOTZ benchmarks are the functions $\{0,1\}^{n}\rightarrow\mathbb{R}^{2}$ defined by
$\textsc{OneMinMax}(x)$ $=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
$\textsc{LOTZ}(x)$ $=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})}\right)$
for all $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
7.2 Runtime
@@
We now prove runtime guarantees for the SMS-EMOA on the OneMinMax and LOTZ benchmarks. Note that the Pareto front, and, more generally, the size of any set of incomparable solutions for OneMinMax and LOTZ have sizes at most $n+1$. Hence we can use Lemma 4 with $\overline{M}=n+1$.
Theorem 26 below gives an upper bound for the runtime of SMS-EMOA on OneMinMax. This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee $\frac{2e^{2}}{e-1}\mu n(\ln n+1)$ for the NSGA-II [ZD23a, Theorem 2]. Due to the more complex population dynamics, the leading constant in the result in [ZD23a] is inferior (58% larger). More critically, the result in [ZD23a] was proven for population sizes $\mu\geq 4(n+1)$ only, and it was disproven for $\mu=n+1$ [ZD23a, Theorem 12], whereas our result here holds for all $\mu\geq n+1$.
@@
Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our result consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
@@
Proof.
Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 4, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations. Assume that for some $i\in[0..n]$, the population $P_{t}$ contains a solution with objective value $(i,n-i)$. Then the probability to reach a solution with value $(i+1,n-i-1)$ is at least
$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$, respectively. Hence, the expected number of iterations to cover the full Pareto front is
$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
We now turn to the LOTZ benchmark. Our runtime guarantee, see Theorem 27 below, is again of the same asymptotic order of magnitude as the guarantee $\frac{2e^{2}}{e-1}\mu n^{2}$ proven for the NSGA-II [ZD23a, Theorem 8], and again, the guarantee for the NSGA-II has a weaker leading constant and requires a larger population size of $\mu\geq 4(n+1)$. In the proof, we first regard the time to reach the individual $1^{n}$, and then consider the times to generate the individuals $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other. In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost.
@@
Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations (or $\mu+2e\mu n^{2}$ fitness evaluations) in expectation, the full Pareto front is covered.
Proof.
Denote by $f$ the LOTZ function. We first consider the time to reach $1^{n}$. Let $X_{t}$ denote the maximum $f_{1}$-value of an individual in $P_{t}$. Assume that at some $t$, we have $X_{t}<n$. Let $x\in P_{t}$ be an individual with $f_{1}(x)=X_{t}$. Then with probability at least $\frac{1}{\mu}(1-\frac{1}{n})^{n-1}\frac{1}{n}$, this $x$ is chosen as parent and mutated to an individual $y$ with $f_{1}(y)>X_{t}$. By Lemma 4, $P_{t+1}$ contains an individual $z$ with $z\succeq y$. In particular, we have $f_{1}(z)\geq f_{1}(y)>X_{t}$ and hence $X_{t+1}>X_{t}$.
We have just seen that the probability to increase $X_{t}$ by at least one is at least $\frac{1}{\mu}(1-\frac{1}{n})^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$. Consequently, the expected waiting time for this event is at most $e\mu n$. After at most $n$ such increases, hence after an expected time of at most $e\mu n^{2}$, we have reached the state $X_{t}=n$, that is, $1^{n}\in P_{t}$.
Once $1^{n}$ is in the population, we consider the time to reach the Pareto optima $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one. Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is in the population. Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ as parent and flip the $i$-th bit from $1$ to 00, with the other bits unchanged. This happens with probability
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
From Lemma 4, we know that any Pareto front point will be maintained in all future generations once reached. Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$. Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$. Since we have at most $n$ such Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
8 Conclusion
@@
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] for the NSGA-III supports our conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
@@
We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.
@@
On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain unchanged when increasing the number of objectives.
@@
Given that our first result shows advantages of the SMS-EMOA and there is only one previous work performing rigorous runtime analyses for this algorithm, we extended our knowledge in this direction by proving competitive runtime guarantees for this algorithm on the two most prominent bi-objective benchmarks. They show that the SMS-EMOA performs here at least at well as the NSGA-II.
Subsequent Work
In [WD24], the runtime of several MOEAs on many-objective problems was analyzed. Concerning the runtime of the classic SMS-EMOA on the $m$OJZJ benchmark, our guarantee of $O(\mu Mn^{k})$ was improved to $O(\mu mn^{k})$. To prove this result, our general technical lemma for the survival of individuals in the population (Lemma 4) was used. The improvement then stemmed from detecting a faster way how all the MOEAs regarded in that work discover the Pareto front. The work [WD24] did not regard the SMS-EMOA with stochastic population update and with heavy-tailed mutation. Since both shorten the time to find an additional extremal block (see the proof of Lemma 8), it is clear that analogous improvements are possible for the SMS-EMOA with stochastic population update and heavy-tailed mutation, giving the bounds $O(\mu mn^{k}\min\{1,\mu/2^{k}\})$ and $O(\mu mk^{\beta-0.5-k}n^{k})$ respectively.